# Timelike conjugate points in Lorentzian length spaces

### v1

@@ 3.1 Symmetric conjugate points
For the reverse implication we use a generalization, due to D. Szeghy [21], of a classical result by F. W. Warner [23]. Let $\gamma(t)=\exp_{p}(tv)$ be a timelike geodesic and let $q=\gamma(1)$ be conjugate to $p$ along $\gamma$. Denoting by $C(p)\subseteq T_{p}M$ the conjugate locus of $p$ we thus have $v\in C(p)$. Let $U\subseteq T_{p}M$ be an open neighborhood of $v$ consisting entirely of timelike vectors. By [21, Thm. 2.1], together with [21, Rem. 4.1], both the set $C^{R}(p)$ of regular conjugate vectors and the set $C(p)^{\prime}$ of conjugate vectors of locally minimal multiplicity are open and dense in $C(p)$, hence so is their intersection $C^{R}(p)\cap C(p)^{\prime}$. Pick $v^{\prime}\in C^{R}(p)\cap C(p)^{\prime}\cap U$. Then by [21, Lem. 4.2], $\exp_{p}$ is a Warner map on a suitable open neighborhood $V\subseteq U$ of $v^{\prime}$ (note that by [21, Sec. 5], the fact that $v^{\prime}$ is timelike implies that $\mathcal{D}_{v^{\prime}}$ is spacelike, so the assumptions of [21, Lem. 4.2] are indeed satisfied). Therefore (cf. [21, Thm. 2.2] and the remarks following it), $\exp_{p}$ is not injective on any neighborhood of $v^{\prime}$. In particular, it is not injective on $U$. Consequently, there exist vectors $v_{i},w_{i}\in U,v_{i}\neq w_{i}$ with $v_{i},w_{i}\to v$ and $\exp_{p}(v_{i})=\exp_{p}(w_{i})\eqqcolon q_{i}$. The distinct timelike geodesics $\gamma_{i}(t)\coloneqq\exp_{p}(tv_{i})$ and $\sigma_{i}(t)\coloneqq\exp_{p}(tw_{i})$ then satisfy one-sided conjugacy. ∎
@@
Clearly, any one-sided conjugate point is also symmetrically conjugate. The converse is open even in the metric case, cf. [20, Open Problem 2.5].
@@ 3.2 Families of timelike geodesics
Observe that since $\ll$ is an open relation (cf. [12, Prop. 2.13]), whenever $x\ll y$ there exist neighborhoods $U$ of $x$ and $V$ of $y$ such that $p^{\prime}\ll q^{\prime}$ for all $p^{\prime}\in U$ and $q^{\prime}\in V$.
@@
Another way to express this is to say that there is exists a unique germ of timelike families around $(\gamma(0),\gamma(1))$ with the respective continuity property. Given a timelike family $F\colon U\times V\to\operatorname{TGeo}(X)$ about $\gamma$, we shall denote by $[F]$ the germ around $(\gamma(0),\gamma(1))$ determined by $F$. Note that if a timelike family is continuous about $\gamma$ and unique among timelike families that are continuous at $\gamma$, then it is also unique among timelike families that are continuous about $\gamma$.
A $C^{0}$-continuous family is a family as above that is continuous w.r.t. $d_{F}$ instead of $d_{\Gamma}$. As we saw before, if $X$ is locally weakly causally closed, localizable, strongly causal and locally maximizing, then these two notions coincide, cf. Proposition 2.26.
@@
Our assumptions mean that (the second alternative in) (i) in Definition 3.7 is satisfied, but (ii) is not. Hence there exists some timelike family $F^{\prime}\colon U^{\prime}\times V^{\prime}\to\operatorname{TGeo}(X)$ continuous at $\gamma$ that does not coincide with $F$ on sufficiently small neighborhoods of $(p,q)$. For each $i\in\mathbb{N}$ pick neighborhoods $U_{i}\subseteq U\cap U^{\prime}$ of $p$ and $V_{i}\subseteq V\cap V^{\prime}$ such that $F|_{U_{i}\times V_{i}}\neq F^{\prime}|_{U_{i}\times V_{i}}$, as well as $U_{i}\subseteq B^{d}_{1/i}(p)$ and $V_{i}\subseteq B^{d}_{1/i}(q)$. Furthermore, by continuity of $F$ and $F^{\prime}$ at $(p,q)$ we may assume $U_{i}$ and $V_{i}$ small enough to grant that $F(U_{i}\times V_{i})\subseteq B^{d_{\Gamma}}_{1/i}(\gamma)$ and $F^{\prime}(U_{i}\times V_{i})\subseteq B^{d_{\Gamma}}_{1/i}(\gamma)$. Pick $(p_{i},q_{i})\in U_{i}\times V_{i}$ such that $\gamma_{i}\coloneqq F(p_{i},q_{i})\neq F^{\prime}(p_{i},q_{i})\eqqcolon\sigma_{i}$. Then $\gamma_{i}$ and $\sigma_{i}$ are timelike geodesics as required to establish that $p$ and $q$ are symmetrically conjugate. ∎
As an immediate consequence we obtain the following analogue of [20, Thm. 3.7]
@@
    For all sequences $p_{i}\to p$ and $q_{i}\to q$ there exists $i_{0}\in\mathbb{N}$ such that for all $i\geq i_{0}$ there exists a geodesic $\gamma_{i}\in\operatorname{TGeo}(X)$ from $p_{i}$ to $q_{i}$ such that $\gamma_{i}\to\gamma$ with respect to $d_{\Gamma}$.
@@
The previous result is a direct analogue of [20, Lem. 3.4]. Note, however, that the embeddability assumption we make is absent from that result, although we are under the impression that it is needed to finish the proof.
@@
Let $(M,g)$ be a strongly causal spacetime. Then any $d_{\Gamma}$-convergent sequence $(\gamma_{i})_{i\in\mathbb{N}}$ of geodesics is strongly embeddable.
@@
Set $p_{i}\coloneqq\gamma_{i}(0)$, $q_{i}\coloneqq\gamma_{i}(1)$, $\gamma\coloneqq\lim_{i}\gamma_{i}$, $p\coloneqq\lim_{i}p_{i}$, $q\coloneqq\lim_{i}q_{i}$ and $v_{i}\coloneqq\gamma_{i}^{\prime}(0)$. By Lemma 2.29, the $v_{i}$ converge to some $v\in T_{p}M$. Choose an open neighborhood $U$ around $p$ that is diffeomorphic to a ball in $\mathbb{R}^{n}$ and let $V$ be some open neighborhood of $q$ in such a way that $U\ll V$. Since $TU\cong U\times\mathbb{R}^{n}\cong\mathbb{R}^{n}\times\mathbb{R}^{n}$ and since $C\coloneqq\{(p_{i},q_{i})\colon i\in\mathbb{N}\}\cup\{(p,q)\}$ is closed, a component-wise application of the Tietze extension theorem provides a continuous map $H\colon U\times V\to TU$ with $H(p_{i},q_{i})=v_{i}$ for $i$ large and $H(p,q)=v$. Compactness of $C$ allows us to additionally shrink $U$ and $V$ to achieve that $H(U\times V)$ lies within the domain of $\exp$ and consists entirely of timelike vectors. Now set $F\colon U\times V\to\operatorname{TGeo}(X)$, $F(x,y)\coloneqq t\mapsto\exp(tH(x,y))$. Then $F$ is continuous, $F(p_{i},q_{i})=\gamma_{i}$ for $i$ large, and $F(p,q)=\gamma$, as required. ∎
@@
Let $(M,g)$ be a strongly causal spacetime and let $\gamma\colon[0,1]\to M$ be a future-directed timelike geodesic connecting $p$ to $q$. Then the following are equivalent:
@@
Applying Proposition 3.15 to the constant sequence $(\gamma)$ it follows that there exists a timelike family that is continuous about $\gamma$. Moreover, Proposition 3.5 shows that (i) is equivalent to $p$ and $q$ not being symmetrically conjugate. The claim therefore follows from Proposition 3.10. ∎
@@
Let $(X,d,\ll,\leq,\tau)$ be a regular, globally hyperbolic and locally maximizing Lorentzian length space. Let $p\ll q$ and suppose there is a unique maximizer $\gamma\in\operatorname{TGeo}(X)$ from $p$ to $q$. Then there exists a timelike family around $\gamma$ consisting of maximizers which is continuous at $\gamma$, in particular, $\gamma$ is embeddable.
@@ 3.3 Unreachable and ultimate conjugate points
Consider n-dimensional anti-de Sitter space $H^{1,n}(K)$ of conjugate radius $\frac{\pi}{\sqrt{-K}}$, $K<0$. Then any two timelike related points $p\ll q$ with $\tau(p,q)=\frac{\pi}{\sqrt{-K}}$ are unreachable conjugate along any timelike maximizer between them. Indeed, any future-directed timelike geodesic $\gamma$ connecting $p$ and $q$ lies in the intersection of $H^{1,n}(K)$ with some plane $P$ connecting $p$, $q$ and 0. Now let $Q\neq P$ be another plane through 0, $p$ and $q$ such that $Q$ cuts out from $H^{1,n}(K)$ a future-directed timelike geodesic $\sigma$ from $p$ to $q$. Then for points $p_{i},q_{i}\in\sigma$ with $p\ll p_{i}$, $q_{i}\ll q$, $p_{i}\to p$ and $q_{i}\to q$, $\sigma$ is the only future-directed timelike geodesic connecting them.
@@
In [20, Ex. 4.7], an example (due to V. Schroeder) of a geodesic $\sigma\colon[0,1]\to N$ in a complete Riemannian manifold $(N,h)$ (specifically, a two-dimensional ellipsoid) is given, with the following properties: $\sigma$ is the unique geodesic in $N$ connecting $x=\sigma(0)$ and $y=\sigma(1)$, while $x$ and $y$ are conjugate along $\sigma$. We may “Lorentzify” this example as follows: Let $M\coloneqq\mathbb{R}\times N$ with Lorentzian metric $g=-dt^{2}+h_{x}$. Then $(M,g)$ is globally hyperbolic (each slice $\{t=t_{0}\}$ is a Cauchy surface), and the following correspondences hold (cf., e.g., [17, Ch. 7]): Any geodesic in $(M,g)$ is of the form $s\mapsto(as+b,\beta)$, where $a,b\in\mathbb{R}$ and $\beta$ is a geodesic in $(N,h)$. Jacobi fields in $(M,g)$ take the form $s\mapsto(T(s),J(s))$, where $T$ is affine linear and $J$ is a Jacobi field in $(N,h)$. Thus if we set $\gamma(s)\coloneqq(as,\sigma(s))$ with $\sigma$ as above, then for $a>0$ sufficiently large, $\gamma$ is a timelike geodesic connecting $p=(0,x)$ to $q=(a,y)$. It is the unique timelike geodesic in $(M,g)$ connecting $p$ and $q$, and these points are conjugate along $\gamma$. Since $\sigma$ is minimizing from $x$ to $y$ in $N$, $\gamma$ is a timelike maximizer from $p$ to $q$ in $(M,g)$. By Corollary 3.22, $p$ and $q$ are reachable along $\gamma$. Consequently, they provide an example of points along a timelike geodesic that are (by Proposition 3.5) symmetrically (hence ultimate) conjugate, but not unreachable conjugate.
@@ 3.4 Causal cut loci
Let $(X,d,\ll,\leq,\tau)$ be a finitely $\tau$-measurable Lorentzian pre-length space and $p\in X$. Given $\gamma\in\operatorname{Geo}(X)$, we say $q\coloneqq\gamma(1)$ is an initial future causal cut point of $p$ along $\gamma$ if
@@
In the case where there is always a $q^{\prime}$ as in (ii), we take $p$ to be the initial future causal cut point of itself along $\gamma$. We denote the set of initial future causal cut points of $p$ along causal geodesics by $\mathrm{IniCut}_{C}^{+}(p)\subseteq X$, called the initial future causal cut locus of $p$. If $X$ is regular, we define the initial future timelike cut locus $\mathrm{IniCut}_{T}^{+}(p)$ and the initial future null cut locus $\mathrm{IniCut}_{N}^{+}(p)$ by replacing all instances of causal with timelike resp. null. It is then clear that $\mathrm{IniCut}_{C}^{+}(p)$ is the disjoint union of $\mathrm{IniCut}_{T}^{+}(p)$ and $\mathrm{IniCut}_{N}^{+}(p)$.
@@
Let $(M,g)$ be a smooth, globally hyperbolic spacetime. Recall that if $\gamma\colon[0,1]\to M$ is a future causal geodesic, then $\gamma(1)$ is in the future causal cut locus (in the standard sense) of $p$ if and only if there exists another maximizer from $p$ to $q$ ($\gamma$ is also maximizing up to $q$) or $q$ is the first conjugate point to $p$ along $\gamma$ (the former case corresponds $q_{i}=q$ in Definition 3.28, see e.g. [22, Prop. 3.2.30]). Hence, $\mathrm{IniCut}_{C}^{+}(p)$ agrees with the usual notion of future causal cut locus in this setting. Similarly, $\mathrm{IniCut}_{T}^{+}(p)$ and $\mathrm{IniCut}_{N}^{+}(p)$ agree with the usual notions of future timelike and future null cut locus, respectively.
@@
In [20, Def. 7.2], the authors require the $q_{i}$ from Definition 3.28 (in their metric setting) to lie on $\gamma$. It is unclear to us whether that definition reduces to the usual notion of cut point in the smooth setting.
@@
Our main result of this section is the following compatibility property.
@@
Let $r\geq 0$ be such that all causally related points of distance at most $r$ can be connected by a unique maximizer, let $p\in X$ and $q\in\mathrm{1stInj}_{C}^{+}(p)$. Then by definition, either there exist two maximizers from $p$ to $q$ (in which case $d(p,q)\geq r$), or, arbitrarily close to $q$ there exist cut points of $p$. In the latter case, if $d(p,q)<r$, let $\varepsilon>0$ and $\tilde{q}$ a cut point of $p$ so that $d(p,\tilde{q})\leq d(p,q)+\varepsilon<r$, which is a contradiction. So in any case, $d(p,q)\geq r$, and taking the infimum over $q$ and then $p$, as well as the supremum over $r\geq 0$, we get
$\mathrm{1stInj}_{C}^{+}(X)\geq\mathrm{UniqueInj}_{C}^{+}(X).$ (3.6)
@@
$\mathrm{1stInj}_{C}^{+}(X)\leq\mathrm{1stInj}_{C}^{+}(p)\leq d(p,q)<R+\varepsilon.$

### v2

@@ 3.1 Symmetric conjugate points
For the reverse implication we use a generalization, due to D. Szeghy [22], of a classical result by F. W. Warner [24]. Let $\gamma(t)=\exp_{p}(tv)$ be a timelike geodesic and let $q=\gamma(1)$ be conjugate to $p$ along $\gamma$. Denoting by $C(p)\subseteq T_{p}M$ the conjugate locus of $p$ we thus have $v\in C(p)$. Let $U\subseteq T_{p}M$ be an open neighborhood of $v$ consisting entirely of timelike vectors. By [22, Thm. 2.1], together with [22, Rem. 4.1], both the set $C^{R}(p)$ of regular conjugate vectors and the set $C(p)^{\prime}$ of conjugate vectors of locally minimal multiplicity are open and dense in $C(p)$, hence so is their intersection $C^{R}(p)\cap C(p)^{\prime}$. Pick $v^{\prime}\in C^{R}(p)\cap C(p)^{\prime}\cap U$. Then by [22, Lem. 4.2], $\exp_{p}$ is a Warner map on a suitable open neighborhood $V\subseteq U$ of $v^{\prime}$ (note that by [22, Sec. 5], the fact that $v^{\prime}$ is timelike implies that $\mathcal{D}_{v^{\prime}}$ is spacelike, so the assumptions of [22, Lem. 4.2] are indeed satisfied). Therefore (cf. [22, Thm. 2.2] and the remarks following it), $\exp_{p}$ is not injective on any neighborhood of $v^{\prime}$. In particular, it is not injective on $U$. Consequently, there exist vectors $v_{i},w_{i}\in U,v_{i}\neq w_{i}$ with $v_{i},w_{i}\to v$ and $\exp_{p}(v_{i})=\exp_{p}(w_{i})\eqqcolon q_{i}$. The distinct timelike geodesics $\gamma_{i}(t)\coloneqq\exp_{p}(tv_{i})$ and $\sigma_{i}(t)\coloneqq\exp_{p}(tw_{i})$ then satisfy one-sided conjugacy. ∎
@@
Clearly, any one-sided conjugate point is also symmetrically conjugate. The converse is open even in the metric case, cf. [21, Open Problem 2.5].
@@ 3.2 Families of timelike geodesics
Observe that since $\ll$ is an open relation (cf. [13, Prop. 2.13]), whenever $x\ll y$ there exist neighborhoods $U$ of $x$ and $V$ of $y$ such that $p^{\prime}\ll q^{\prime}$ for all $p^{\prime}\in U$ and $q^{\prime}\in V$.
@@
Another way to express this is to say that there exists a unique germ of timelike families around $(\gamma(0),\gamma(1))$ with the respective continuity property. Given a timelike family $F\colon U\times V\to\operatorname{TGeo}(X)$ about $\gamma$, we shall denote by $[F]$ the germ around $(\gamma(0),\gamma(1))$ determined by $F$. Note that if a timelike family is continuous about $\gamma$ and unique among timelike families that are continuous at $\gamma$, then it is also unique among timelike families that are continuous about $\gamma$.
A $C^{0}$-continuous family is a family as above that is continuous w.r.t. $d_{F}$ instead of $d_{\Gamma}$. As we saw before, if $X$ is localizable, locally weakly causally closed, strongly causal and locally maximizing, then these two notions coincide, cf. Proposition 2.26.
@@
Our assumptions mean that (the second alternative in) (i) in Definition 3.7 is satisfied, but (ii) is not. Hence there exists some timelike family $F^{\prime}\colon U^{\prime}\times V^{\prime}\to\operatorname{TGeo}(X)$ continuous at $\gamma$ that does not coincide with $F$ on sufficiently small neighborhoods of $(p,q)$. For each $i\in\mathbb{N}$ pick neighborhoods $U_{i}\subseteq U\cap U^{\prime}$ of $p$ and $V_{i}\subseteq V\cap V^{\prime}$ such that $F|_{U_{i}\times V_{i}}\neq F^{\prime}|_{U_{i}\times V_{i}}$, as well as $U_{i}\subseteq B^{d}_{1/i}(p)$ and $V_{i}\subseteq B^{d}_{1/i}(q)$. Furthermore, by continuity of $F$ and $F^{\prime}$ at $(p,q)$ we may assume $U_{i}$ and $V_{i}$ small enough to grant that $F(U_{i}\times V_{i})\subseteq B^{d_{\Gamma}}_{1/i}(\gamma)$ and also $F^{\prime}(U_{i}\times V_{i})\subseteq B^{d_{\Gamma}}_{1/i}(\gamma)$. Pick $(p_{i},q_{i})\in U_{i}\times V_{i}$ such that $\gamma_{i}\coloneqq F(p_{i},q_{i})\neq F^{\prime}(p_{i},q_{i})\eqqcolon\sigma_{i}$. Then $\gamma_{i}$ and $\sigma_{i}$ are timelike geodesics as required to establish that $p$ and $q$ are symmetrically conjugate. ∎
As an immediate consequence we obtain the following analogue of [21, Thm. 3.7]
@@
    For all sequences $p_{i}\to p$ and $q_{i}\to q$ there exists $i_{0}\in\mathbb{N}$ such that for all $i\geq i_{0}$ there is a geodesic $\gamma_{i}\in\operatorname{TGeo}(X)$ from $p_{i}$ to $q_{i}$ such that $\gamma_{i}\to\gamma$ with respect to $d_{\Gamma}$.
@@
The previous result is a direct analogue of [21, Lem. 3.4]. Note, however, that the embeddability assumption we make is absent from that result, although we are under the impression that it is needed to finish the proof.
@@
Let $(M,g)$ be a strongly causal smooth spacetime. Let $(\gamma_{i})_{i\in\mathbb{N}}$ be a sequence of timelike geodesics in $M$ converging (with respect to $d_{\Gamma}$) to a limit geodesic $\gamma$. Let $p=\gamma(0)$ and $q=\gamma(1)$. If $p$ and $q$ are not conjugate along $\gamma$, then the sequence $(\gamma_{i})$ is strongly embeddable.
@@
Let $v_{i}=\gamma_{i}^{\prime}(0)\in TM$ and $v=\gamma^{\prime}(0)\in T_{p}M$. By Lemma 2.29, we have $v_{i}\to v$. Consider the map $E:TM\supseteq\mathcal{D}\to M\times M$ defined on a suitable open neighbourhood $\mathcal{D}$ of the 0-section in TMTM by
$E(w)=(\pi(w),\exp_{\pi(w)}(w)),$
with $\pi:TM\to M$ the bundle projection. Then the differential of $E$ at $v$ takes the following lower-triangular block form:
$T_{v}E\cong\begin{pmatrix}\mathrm{id}&0\\ \star&T_{v}\exp_{p}\end{pmatrix}$.
As $p$ and $q$ are not conjugate along $\gamma$, $T_{v}\exp_{p}$ is invertible, implying that $E$ is a diffeomorphism from a neighbourhood $W\subset TM$ of $v$ onto a neighbourhood $U\times V$ of $(p,q)$ in $M\times M$. Let $H:U\times V\to W$ be the smooth inverse of $E|_{W}$.
We define the timelike family $F:U\times V\to\operatorname{TGeo}(M)$ by
$F(x,y)(t):=\exp_{x}(t\cdot H(x,y)).$
This family is continuous about $\gamma$. Moreover, since $\gamma_{i}\to\gamma$, for sufficiently large $i$ the initial velocities $v_{i}$ lie in $W$ and the endpoints $(p_{i},q_{i})$ lie in $U\times V$. Consequently, $H(p_{i},q_{i})=v_{i}$, and $F(p_{i},q_{i})$ is exactly the geodesic $\gamma_{i}$. Thus, the sequence is strongly embeddable. ∎
The following result is a direct analogue of [21, Prop. 3.5].
@@
Let $(M,g)$ be a smooth strongly causal spacetime and let $\gamma\colon[0,1]\to M$ be a future-directed timelike geodesic connecting $p$ to $q$. Then the following are equivalent:
@@
[…]
@@
Let $(X,d,\ll,\leq,\tau)$ be a regular, globally hyperbolic and locally maximizing Lorentzian length space. Let $p\ll q$ and suppose there is a unique maximizer $\gamma\in\operatorname{TGeo}(X)$ from $p$ to $q$. Then there exists a timelike family about $\gamma$ consisting of maximizers which is continuous at $\gamma$, in particular, $\gamma$ is embeddable.
@@ 3.3 Unreachable and ultimate conjugate points
Consider $n$-dimensional anti-de Sitter space $H^{1,n}(K)$ of conjugate radius $\frac{\pi}{\sqrt{-K}}$, $K<0$. Then any two timelike related points $p\ll q$ with $\tau(p,q)=\frac{\pi}{\sqrt{-K}}$ are unreachable conjugate along any timelike maximizer between them. Indeed, any future-directed timelike geodesic $\gamma$ containing $p$ and $q$ lies in the intersection of $H^{1,n}(K)$ with some plane $P$ connecting $p$, $q$ and 0. Now let $Q\neq P$ be another plane through 0, $p$ and $q$ such that $Q$ cuts out from $H^{1,n}(K)$ a future-directed timelike geodesic $\sigma$ from $p$ to $q$. Then for points $p_{i},q_{i}\in\sigma$ with $p\ll p_{i}$, $q_{i}\ll q$, $p_{i}\to p$ and $q_{i}\to q$, $\sigma$ is the only future-directed timelike geodesic connecting them.
@@
In [21, Ex. 4.7], an example (due to V. Schroeder) of a geodesic $\sigma\colon[0,1]\to N$ in a complete Riemannian manifold $(N,h)$ (specifically, a two-dimensional ellipsoid) is given, with the following properties: $\sigma$ is the unique geodesic in NN connecting $x=\sigma(0)$ and $y=\sigma(1)$, while $x$ and $y$ are conjugate along $\sigma$. We may “Lorentzify” this example as follows: Let $M\coloneqq\mathbb{R}\times N$ with Lorentzian metric $g=-dt^{2}+h_{x}$. Then $(M,g)$ is globally hyperbolic (each slice $\{t=t_{0}\}$ is a Cauchy surface), and the following correspondences hold (cf., e.g., [18, Ch. 7]): Any geodesic in $(M,g)$ is of the form $s\mapsto(as+b,\beta)$, where $a,b\in\mathbb{R}$ and $\beta$ is a geodesic in $(N,h)$. Jacobi fields in $(M,g)$ take the form $s\mapsto(T(s),J(s))$, where $T$ is affine linear and $J$ is a Jacobi field in $(N,h)$. Thus if we set $\gamma(s)\coloneqq(as,\sigma(s))$ with $\sigma$ as above, then for $a>0$ sufficiently large, $\gamma$ is a timelike geodesic connecting $p=(0,x)$ to $q=(a,y)$. It is the unique timelike geodesic in $(M,g)$ connecting $p$ and $q$, and these points are conjugate along $\gamma$. Since $\sigma$ is minimizing from $x$ to $y$ in NN, $\gamma$ is a timelike maximizer from $p$ to $q$ in $(M,g)$. By Corollary 3.22, $p$ and $q$ are reachable along $\gamma$. Consequently, they provide an example of points along a timelike geodesic that are (by Proposition 3.5) symmetrically (hence ultimate) conjugate, but not unreachable conjugate.
@@ 3.4 Causal cut loci
Let $(X,d,\ll,\leq,\tau)$ be a finitely $\tau$-measurable Lorentzian pre-length space and $p\in X$. Given $\gamma\in\operatorname{Geo}(X)$, we say $q\coloneqq\gamma(1)$ is an initial future causal cut point of $p=\gamma(0)$ along $\gamma$ if
@@
In the case where there is always a $q^{\prime}$ as in (ii), we take $p$ to be the initial future causal cut point of itself along $\gamma$. We denote the set of initial future causal cut points of $p$ along causal geodesics by $\mathrm{IniCut}_{C}^{+}(p)\subseteq X$, called the initial future causal cut locus of pp. If $X$ is regular, we define the initial future timelike cut locus $\mathrm{IniCut}_{T}^{+}(p)$ and the initial future null cut locus $\mathrm{IniCut}_{N}^{+}(p)$ by replacing all instances of causal with timelike resp. null. It is then clear that $\mathrm{IniCut}_{C}^{+}(p)$ is the disjoint union of $\mathrm{IniCut}_{T}^{+}(p)$ and $\mathrm{IniCut}_{N}^{+}(p)$.
@@
Let $(M,g)$ be a smooth, globally hyperbolic spacetime. Recall that if $\gamma\colon[0,1]\to M$ is a future causal geodesic, then $\gamma(1)$ is in the future causal cut locus (in the standard sense) of $p$ if and only if there exists another maximizer from $p$ to $q$ ($\gamma$ is also maximizing up to $q$) or $q$ is the first conjugate point to $p$ along $\gamma$ (the former case corresponds $q_{i}=q$ in Definition 3.28, see e.g. [23, Prop. 3.2.30]). Hence, $\mathrm{IniCut}_{C}^{+}(p)$ agrees with the usual notion of future causal cut locus in this setting. Similarly, $\mathrm{IniCut}_{T}^{+}(p)$ and $\mathrm{IniCut}_{N}^{+}(p)$ agree with the usual notions of future timelike and future null cut locus, respectively.
@@
In [21, Def. 7.2], the authors require the $q_{i}$ from Definition 3.28 (in their metric setting) to lie on $\gamma$. This notion, however, does not coincide with the standard notion on smooth Riemannian manifolds. Indeed, the two-dimensional torus of revolution provides a counterexample: The cut locus of a point $p$ slightly above the outer equator contains an arc lying on a parallel whose endpoints are conjugate but cannot be reached by more than one minimizer from $p$, see [10], (iii) on p. 104 and Fig. 4. Consider one of these (two) points, call it $q$ and suppose it is reached from $p$ by a geodesic $\gamma$. Since $\gamma$ beyond $q$ cannot coincide with the arc it does not contain any meeting points of minimizers from $p$ (although $q$ is a limit of such points lying on the arc).
@@
Our main result of this subsection is the following compatibility property.
@@
Let $r\geq 0$ be such that all causally related points of distance at most $r$ can be connected by a unique maximizer, let $p\in X$ and $q\in\mathrm{IniCut}_{C}^{+}(p)$. Then by definition, either there exist two maximizers from $p$ to $q$ (in which case $d(p,q)\geq r$), or, arbitrarily close to $q$ there exist cut points of pp. In the latter case, if $d(p,q)<r$, let $\varepsilon>0$ and $\tilde{q}$ a cut point of $p$ so that $d(p,\tilde{q})\leq d(p,q)+\varepsilon<r$, which is a contradiction. So in any case, $d(p,q)\geq r$, and taking the infimum over $q$ and then $p$, as well as the supremum over $r\geq 0$, we get
$\mathrm{IniInj}_{C}^{+}(X)\geq\mathrm{UniqueInj}_{C}^{+}(X).$ (3.6)
@@
$\mathrm{IniInj}_{C}^{+}(X)\leq\mathrm{IniInj}_{C}^{+}(p)\leq d(p,q)<R+\varepsilon.$